OpenStudy (anonymous):

use the formal definition of limit to verify the indicated limit: limit x-->infinite (1/sqrt(x))

OpenStudy (zarkon):

Pick \[N=\frac{1}{\epsilon^2}\]

OpenStudy (anonymous):

oops, limit x-->infinite (1/sqrt(x)) = 0

OpenStudy (zarkon):

Obviously ;)

OpenStudy (anonymous):

let \[\epsilon\] be a given positive number. for x > 0 we have | 1/sqrt(x) - 0 | = 1/|sqrt(x)|, 1/x < \[\epsilon^2\] provided x>1/\[\epsilon^2\] R = 1/e^2 => R^(1/2)=1/e ???

OpenStudy (anonymous):

is it wrong? :(?

OpenStudy (zarkon):

yes.. \[\text{Let }\epsilon>0\] \[\text{Choose }N=\frac{1}{\epsilon^2}\] then for all \[x>N\] we have \[x>\frac{1}{\epsilon^2}\] \[\Rightarrow \epsilon^2>\frac{1}{x}\] \[\Rightarrow \epsilon>\frac{1}{\sqrt{x}}\]